Lewis Securities Inc. has decided to acquire a new market data and quotation system for its Richmond home office. The system receives current market prices and other information from several online data services and then either displays the information on a screen or stores it for later retrieval by the firm’s brokers. The system also permits customers to call up current quotes on terminals in the lobby. The equipment costs $1,000,000 and, if it were purchased, Lewis could obtain a term loan for the full purchase price at a 10% interest rate. Although the equipment has a 6-year useful life, it is classified as a special-purpose computer and therefore falls into the MACRS 3-year class. If the system were purchased, a 4-year maintenance contract could be obtained at a cost of $20,000 per year, payable at the beginning of each year. The equipment would be sold after 4 years, and the best estimate of its residual value is $200,000. However, because real-time display system technology is changing rapidly, the actual residual value is uncertain. As an alternative to the borrow-and-buy plan, the equipment manufacturer informed Lewis that Consolidated Leasing would be willing to write a 4-year guideline lease on the equipment, including maintenance, for payments of $260,000 at the beginning of each year. Lewis’s marginal federal-plus-state tax rate is 40%. You have been asked to analyze the lease-versus-purchase decision and, in the process, to answer the following questions: 5. Now assume that the equipment’s residual value could be as low as $0 or as high as $400,000, but $200,000 is the expected value. Because the residual value is riskier than the other relevant cash flows, this differential risk should be incorporated into the analysis. Describe how this could be accomplished. (No calculations are necessary, but explain how you would modify the analysis if calculations were required.) What effect would the residual value’s increased uncertainty have on Lewis’ lease-versus-purchase decision? 6. The lessee compares the present value of owning the equipment with the present value of leasing it. Now put yourself in the lessor’s shoes. In a few sentences, how should you analyze the decision to write or not to write the lease? Using complete sentences and academic vocabulary, please answer questions 5 and 6. Please don't copy from other sources. Thank you
Answer #1
5. We can use the most probable or expected cash flow --ie. (50%*0)+(50%*400000)= $ 200000 as the salvage /residual value cash flow in Year 4
Present Value of the Purchase option using the most expected valueof $ 200000 as salvage
Year 0 1 2 3 4
Initial investment 1000000
Annual Maintenance 20000 20000 20000 20000
Depn. Tax shields -133320 -177800 -59240 -29640
After-tax salvage(200000*(1-40%)) -120000
Total annual cash outflow 1020000 -113320 -157800 -39240 -149640
PV F at 10% 1 0.90909 0.82645 0.75131 0.68301
PV at 10% 1020000 -103018 -130413 -29482 -102206
NPV of costs 654881
PV of leasing
with beginning-of-year after-tax lease payments of
260000*(1-40%)=
156000
PV of costs to Lease=
(156000*(1-1.1^-4)/0.1)*(1.1)=
543949

From lessee's point of view,

PV of leasing is LESSER & so recommended.

Year MACRS % Depn. Amt. Tax shield=Depn.amt.*40%
1 33.33 333300 133320
2 44.45 444500 177800
3 14.81 148100 59240
4 7.41 74100 29640
100 1000000 400000
6.The lessor studies the lessee's credit-worthiness to pay the regular lease amounts.
He considers his after-tax benefits out of the lease rental incomes.
Availability of depreciation tax -shields